### v5

@@ 2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness:
@@
where the proportion $P_{ii}$ is generally larger than .7.7.7.7. This illustrates $\pi$’s Matthew effect of accumulated advantage, often summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies exhibit substantial competitiveness compared to less powerful ones, revealing a quadratic relationship between $\pi_{i}$ and $g_{i}$. For instance, the USA’s $\pi_{i}$ could be significantly higher than Japan’s, relative to their production ratio. The amplification also stems from the USA’s relative advantages over other countries compared to Japan’s. In Eq. (7), with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$, respectively,
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$. Alternatively, the Matthew effect is largely explained by preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among countries according to their existing competitiveness.
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. Thus, it exhibits quadratic growth when $\tilde{g}_{i}$ increases linearly, and in a linear regression:
@@
for some $c_{1}>0,c_{2}>0$, where $\epsilon_{i}$ is the residual term. Therefore, when dealing with a rapidly thriving country $i$, country $j$ should mitigate its conflicts with $i$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say:
@@
for some $c_{4}>0$, then $\tilde{g}_{i}$ would be accelerated quadratically, leading to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $\tilde{g}_{i}$ would grow exponentially to their maximum capacities. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
Additionally, the welfare associated with competitive advantage leads to an agglomeration effect, clustering all countries around a select few of economic superpowers. According to Corollary 1, imports from a large economy are preferred over imports from a smaller economy, all else being equal. Consequently, the smaller economy should counter this preference by lowering its export prices. Meanwhile, a third party could impose different tariffs on large and small countries to ensure that the same imported products have uniform prices in the third country. Furthermore, the large market, characterized by numerous producers and consumers where no single entity can dominate, may be favored by the third party for its market competitiveness. In either scenario, a small country tends to fear the emergence of a neighboring superpower that shares similar location advantages and resource endowments.
@@
Assume the Matthew effect. If $g_{j}>g_{k}$, then the third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
Thus, providing additional welfare to competitive economies is another significant force driving trade flows across the network, alongside comparative advantage. However, trade wars and protectionism are more likely due to competitive advantages rather than comparative advantages, which benefit all trade partners. Accordingly, an economic superpower would exert all efforts to maintain its status when facing challenges from emerging competitors; preserving its advantage in $\pi$ prevents the domino effect, which is contrary to the Matthew effect. As Copeland (2014) states, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. An emerging superpower’s aspiration to assume its rightful place would encounter significant challenges. World War I and II were intense conflicts between existing and emerging superpowers, leading to the establishment of new world orders.
@@ 3 Bilateral Trade War for Competitiveness
Anticipating a significant positive change to $\pi_{i}$, country $i$ considers initiating a trade war against its trade partner $j$. Gaming on the matrix $P$, country $i$’s first move could alter the element $P_{ji}$ — the portion of $j$’s production exported to $i$. For instance, raising tariffs or decreasing import quotas on $j$ reduces imports from $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume the bilateral conflict involves only the two counterparties in $P$ without directly affecting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. However, changes in $P_{ji}$ or $P_{ij}$ eventually impact global value chains in multiple steps, indirectly involving all third parties. This simplification may not reflect the complexity of the 2018 China-USA trade war, where the Biden administration urged allies to ban Chinese companies from purchasing advanced chips and chip-making equipment (New York Times, 2022).
@@
In retaliation, country $j$ might decrease $P_{ij}$ or $P_{ji}$ sooner or later. Without retaliation, the war ends at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages but might also invite additional sanctions from $i$ if $j$ shows weakness. Overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ may also severely harm $\pi_{j}$. For example, after the Trump administration imposed tariffs and trade barriers on China in January 2018, China responded in April 2018 by imposing tariffs on 128 products imported from the USA (Washington Post, 2018).
@@
We introduce a matrix $\Lambda$ for retaliation actions. In retaliation for a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s change $\Delta P_{ij}$ in $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. Thus, $\lambda_{ij}=1/\lambda_{ji}$, in the veil of any first-mover advantage. We assume $\lambda_{ji}>0$ to align actions from both countries. Ignoring temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; thus, the reduced exports will be consumed domestically, and the modified $P$ remains a stochastic matrix.
@@
The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. Generally, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. In the following exposition, we first assume the coefficient is an exogenous or predetermined constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
@@
We introduce a few notations for the next five theorems and their corollaries. Let $I_{n}$ be the $n\times n$ identity matrix and let $\pi_{{}_{-i}}$ be the transpose of $\pi$ with $\pi_{i}$ removed. The square matrix $Z_{i}$ is the transpose of $P$ with its $i$th row and $i$th column removed. Also, the column vector $\alpha_{i}$ extracts the $i$th row from $P$ and then drops its $i$th element. As usual, $\vec{1}_{n}$ and $\vec{0}_{n}$ are the $n\times 1$ column vectors with all ones and zeros, respectively. For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has one for its $j$th element and zeros elsewhere. Finally, the vector $\gamma_{ji}$ takes the values of $e_{j}$ and then removes its $i$th element. So, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
@@
With the above setting and notations, for any $j\neq i$,
@@
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain:
$\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
Using Theorem 2, we have the following Corollary 2, already implied in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
@@
When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
Additionally, since
$(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3. In particular, when $\lambda_{ji}=0$ in Eqs. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
@@
By varying $j$ in Eq. (9), we can identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change in $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the sign of $\mathrm{d}P_{ji}$.
In search of the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is

### v6

@@ 2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$’s competitiveness $\pi_{i}$ accumulates in proportion to its initial level of competitiveness:
@@
where the self-reinforcing proportion $P_{ii}$ is typically greater than .7.7. This formulation illustrates the Matthew effect — the phenomenon of accumulated advantage — often summarized by the aphorism: “the rich get richer and the poor get poorer.”
For example, the United States may exhibit a substantially higher $\pi_{i}$ than Japan, compared to their production ratio. This amplification arises not only from the U.S.’s internal advantages but also from its relative superiority over other countries compared to Japan. Specifically, in Eq. (7), for $i=\mathrm{USA}$ or $i=\mathrm{Japan}$:
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}}.\end{array}\right$.
The Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $1>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>0$; and $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$.
Alternatively, the Matthew effect can be interpreted through preferential attachment in endogenous weighting, where weights are distributed proportionally to existing competitiveness. This implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. Normalizing $g_{i}$ as $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{g_{i}}{\sum_{j=1}^{n}g_{j}}$, the effect suggests that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$, after controlling for $\tilde{g}_{i}$. Thus, $\pi_{i}$ exhibits quadratic growth as $\tilde{g}_{i}$ increases linearly. In a linear regression framework:
@@
for some $c_{1}>0,c_{2}>0$, and $\epsilon_{i}$ is the residual term. Therefore, when country $i$ is rapidly ascending, country $j$ should seek to resolve conflicts with $i$ before $i$ becomes too powerful to compromise.
If country $i$ leverages its accelerating $\pi_{i}$ to boost production $g_{i}$, such as:
@@
with $c_{4}>0$, then $\tilde{g}_{i}$ accelerates quadratically, leading to quartic growth in $\pi_{i}$. Iterating this process results in exponential growth of both $\pi_{i}$ and $\tilde{g}_{i}$ toward their maximum capacities. Historically, many empires have emerged rapidly, gaining prominence with astonishing speed.
Moreover, the welfare associated with competitive advantage induces an agglomeration effect, clustering countries around a few economic superpowers. According to Corollary 1, imports from a large economy are preferred over those from a smaller one, all else being equal. Consequently, smaller economies may counteract this preference by lowering export prices. A third party might impose differentiated tariffs on large and small countries to equalize import prices. Additionally, large markets — characterized by numerous producers and consumers without a dominant entity — may be favored for their market competitiveness. In either case, smaller countries often fear the rise of neighboring superpowers with similar geographic and resource advantages.
@@
Assume the Matthew effect. If $g_{j}$ is significantly greater than $g_{k}$, then third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
Thus, the additional welfare granted to competitive economies becomes a major force driving trade flows, alongside comparative advantage. However, trade wars and protectionism are more likely to stem from competitive advantages, which benefit only select partners, rather than comparative advantages, which benefit all. Accordingly, an economic superpower will strive to preserve its status when challenged by emerging competitors. Maintaining its advantage in $\pi$ helps prevent a reverse Matthew effect. As Copeland (2014) notes: “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, the incentive is more proportional to $\pi_{i}$ than to $g_{i}$. Emerging superpowers seeking their rightful place face formidable challenges. World Wars I and II were intense confrontations between established and rising powers, ultimately reshaping the global order.
@@ 3 Bilateral Trade War for Competitiveness
Anticipating a significant positive shift in its competitiveness $\pi_{i}$, country $i$ may consider initiating a trade war against its trade partner $j$. By strategically manipulating the trade matrix $P$, country $i$’s initial move could alter the element $P_{ji}$ — the share of country $j$’s production exported to country $i$. For example, increasing tariffs or reducing import quotas on goods from country $j$ would lower imports from $j$, thereby decreasing $P_{ji}$. Country $i$ might also restrict certain exports to $j$, directly reducing $P_{ij}$.
@@
For simplicity, we assume the bilateral conflict affects only the two counterparties in matrix $P$, without directly impacting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. However, changes in $P_{ji}$ or $P_{ij}$ can propagate through global value chains, indirectly involving third parties over multiple steps. This simplification does not fully capture the complexity of events such as the 2018 China–USA trade war, during which the Biden administration encouraged allies to prohibit Chinese firms from acquiring advanced chips and chip-making equipment (New York Times, 2022).
3.1 Reciprocal Retaliations
@@
In response, country $j$ may retaliate by reducing $P_{ij}$ or $P_{ji}$. For instance, following the Trump administration’s imposition of tariffs and trade barriers on China in January 2018, China retaliated in April 2018 by levying tariffs on 128 U.S. products (Washington Post, 2018a). If country $j$ does not retaliate, the conflict ends at its inception, and country $i$ emerges victorious. A lack of response or unconditional surrender may prevent further damage but could also invite additional sanctions from country $i$ if country $j$ appears weak. Conversely, an overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ could also severely harm $\pi_{j}$.
To formalize retaliation, we introduce a matrix $\Lambda$ representing retaliation actions. In response to a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ adjusts $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for a change $\Delta P_{ij}$, country $j$ modifies $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. Thus, $\lambda_{ij}=1/\lambda_{ji}$, assuming no first-mover advantage.
@@
We assume $\lambda_{ji}>0$ to ensure aligned actions from both countries. Ignoring temporally lagged reactions on the long-run effects $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ yields equivalent impacts. Letting $\lambda_{ii}=1$, we compile all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, for all $i,j\in\mathcal{N}$. To preserve the unit sum in each row of $P$, we deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$. The reduced exports are then consumed domestically, ensuring that the modified $P$ remains a stochastic matrix.
@@
The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both countries, including political considerations, anti-dumping measures, trade deficits, and the nature of traded commodities or services. In practice, there is no universal formula for every contingency, and negotiations may involve multiple rounds of bilateral discussions. In the subsequent section, we assume $\lambda_{ji}$ is an exogenous or predetermined constant in a potential trade war. Section 4 explores a generic yet simple bargaining solution for $\lambda_{ji}$ across most of the n(n1)/2n(n-1)/2 bilateral conflicts. The solution aims not to eliminate conflicts but to mitigate them and fairly distribute the associated costs, serving the diverse interests of the involved countries. In a mixed cooperative and noncooperative setting, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ may be of limited interest to either party.
@@
We introduce several notations for the next five theorems and their corollaries. Let $I_{n}$ denote the $n\times n$ identity matrix and $\pi_{{}_{-i}}$ represent the transpose of $\pi$ with its $i$th element removed. Define $Z_{i}$ as the transpose of $P$ with its $i$th row and column removed. The column vector $\alpha_{i}$ is obtained by extracting the $i$th row from $P$ and removing its $i$th element. Vectors $\vec{1}_{n}$ and $\vec{0}_{n}$ are $n\times 1$ column vectors of all ones and all zeros, respectively. For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has a $1$ in its $j$th position and zeros elsewhere. Finally, the vector $\gamma_{ji}$ is derived from $e_{j}$ by removing its $i$th element. Thus, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
@@
Under the above setting and notations, for any $j\neq i$:
@@
Using the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also have:
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\qquad\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\qquad\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}=\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.$
Theorem 2 implies Corollary 2, a variant of Eq. (6) — $\pi_{i}$ remains unchanged when the authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) does not rely on the assumptions in Theorem 2.
@@
When $\lambda_{ji}=\pi_{j}/\pi_{i}$ (or equivalently, $\lambda_{ij}=\pi_{i}/\pi_{j}$), $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}=\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}=\vec{0}_{n}^{\top}$.
Additionally, since $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ has all non-negative elements, the derivatives $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3. In a trade friction (i.e., $\mathrm{d}P_{ji}<0$), $j$’s overreaction with $\lambda_{ji}>\pi_{j}/\pi_{i}$ benefits $\pi_{i}$ but deteriorates $\pi_{j}$, according to Eqs. (9)–(10).
@@
3.2 Identification of Competitors and Collaborators
By varying country $j$ in Eq. (9), we can identify country $i$’s trade collaborators and competitors. In Eq. (9), the derivative $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has the opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. Accordingly, if $\lambda_{ji}>\pi_{j}/\pi_{i}$, then country $j$ becomes a potential target for conflict with country $i$ because a small negative change in $P_{ji}$ would result in a positive change in $\pi_{i}$. Conversely, if $\lambda_{ji}<\pi_{j}/\pi_{i}$, then country $j$ is a potential candidate for deeper collaboration. In this case, when $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, a larger $\lambda_{ji}$ slows the growth of $\pi_{i}$ and accelerates the decline of $\pi_{j}$. Finally, if country $j$ maintains $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any small change in $P_{ji}$ will have negligible effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the direction of $\mathrm{d}P_{ji}$.
To facilitate cross-country comparisons, we express changes in terms of percentage variation in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that in $P_{ji}$ is given by: